### v2

@@ 5.2 Approximate analytical expressions for the two-point function
where $R$ is the radius of the semicircle distribution. For the RP model, we find that in the regime of large $N$ and small $\gamma$, the radius scales approximately as $R=2/N^{\gamma/2}$, consistent with the structure of the model’s Hamiltonian [^13] For $N=500$, we observed that the Wigner semicircle described well the DOS as long as $\gamma\leq 0.5$..
Substituting Eq. (52) into Eq. (51), we obtain an analytical approximation for the two-point function:
$F_{2}^{(\gamma)}(t)\approx\left(\frac{2J_{1}(Rt)}{Rt}\right)^{2}\,,\qquad R=\frac{2}{N^{\gamma/2}}\,,$ (53)
where $J_{1}$ denotes the Bessel function of the first kind. For $N=500$, we find that this expression accurately captures the early-time behavior of the two-point function for $\gamma\lesssim 0.5$. For $\gamma=0$ (53) reduces to the result derived in Cotler:2017jue :
$F_{2}^{(\gamma=0)}(t)=\frac{J_{1}(2t)^{2}}{t^{2}}\,.$ (54)
Figure 5 (left and middle) shows the numerical results of the $2$-point and $4$-point functions for $L=8$. The analytic expression (53) matches with the numerical results at early times for small values of $\gamma$. These cases are dominated by the off-diagonal part of the GOE Hamiltonian (20). The results obtained with (53) are shown by the black dashed line in the left panel of Fig. 5. As $\gamma$ increases, the decay develops a longer tail. On the other hand, the 4-point functions exhibit clear zeroes and sharp local minima in the time axis for all $\gamma$. The decay is also faster than that of the 2-point functions. To analyze this closely, we study the behavior of $2n$-point functions for $n=1,\cdots,5$ with $\gamma=0$ for a system size of $L=8$. This is illustrated in the rightmost panel of Fig. 5. As observed, for a fixed $\gamma$, higher-point functions exhibit a faster decay.
@@ 5.3 Cumulative OTOCs
Figures 6 and 7 illustrate the behaviour of the cumulative partial sum $\mathcal{F}_{p}^{(\gamma)}(t):=\mathcal{F}_{p}^{(\gamma)}(f\,;t)$ for the choice of weights $f(n)=1/n$ and $f(n)=1/n^{2}$ respectively in the ergodic (left) and fractal (right) regimes, for $s=1/2$ and $L=8$. Since higher-point functions decay faster than lower-point functions Cotler:2017jue , in practice, only a finite $p$ suffices to produce reasonable results.
@@
$\sum_{n=1}^{\infty}\frac{1}{n^{k}}e^{-2n\chi t}=\textrm{Li}_{k}(e^{-2\chi t})\approx e^{-2\chi t}\,\quad(\mathrm{large}~{}t)\,.$ (59)
This is indeed representative of the numerical results for the ergodic phase (left) in Figs. 6 and 7, where the horizontal red dashed line in the latter case corresponds precisely to $\zeta(2)=\pi^{2}/6$. The exponential decay is represented by the black dashed line in Figs. 6 and 7 (left), demonstrating the eventual vanishing of the cumulative OTOC. The $\gamma=0$ result primarily shows the result when the Hamiltonian is strictly controlled by the GOE matrices in the RP Hamiltonian. The same conclusion holds for other classes of Gaussian ensembles.
@@
In the fractal phase, the cumulative sum undergoes a decay, albeit at a much slower rate compared to the ergodic regime. As illustrated in the inset of Figs. 6 and 7 (right), at late times, the cumulative OTOC does not approach zero; rather, it decays to a small, non-zero value. This phenomenon reflects a form of “memory” retained within the operator statistics, which we will see in the subsequent section. In contrast, within the localized phase, the cumulative partial sum exhibits minimal decay, remaining almost constant at its initial value.
6 Operator statistics from free probability
In this section, we investigate the statistical properties of operators within the framework of the RP model. Specifically, our focus lies on the eigenvalue statistics of the following operator:
@@ 6 Operator statistics from free probability
where the temporal evolution is generated by the RP model Hamiltonian, as defined in (20). Here $S_{1}(0)$ and $S_{L}(0)$ represent the spin operators at the initial time $t=0$, defined in (46) and (47). The total sum of the operators $S(t)$ constructed through this formalism, encapsulates the impact of the RP Hamiltonian (20) over time. Our primary objective is to systematically analyze the eigenvalue statistics of $S(t)$ at various time instants $t$. By varying $\gamma$, i.e., the parametric regime of the RP Hamiltonian, distinct phases (ergodic, fractal, or localized) or structures in the operator statistics are anticipated, potentially exhibiting novel characteristics inspired by predictions from free probability theory.
@@ 6.1 Choice of the initial operators
In the following analysis, we consider the operators $S_{1}(0)$ and $S_{L}(0)$ to be spin operators associated with spin $s$. This formulation generalizes the construction and enables comparisons against free probability predictions for various spin values, as obtained through analytical methods outlined in Section 2. The operators $S_{1}(0)$ and $S_{L}(0)$ are defined as tensor products of spin operators as in (46)-(47).
Figure 8: The statistics of the eigenvalues of the time-evolved $s=1/2$ operator (60) at the initial time $t=0$ (top, left) and late times $t=50$ in the ergodic (top, right), fractal (bottom, left), and localized (bottom, right) regimes. We choose $L=8$ (i.e., dimension of the Hamiltonian $N=(2s+1)^{L}=256$) with $5000$ ensemble averages. The solid blue lines describe the arsine distribution (12) i.e., Fig. 1 (a).
@@ 1 $S_{1},S_{L}$$\{1,0,-1\}$
This is a convention that we follow throughout our discussion. More generally, we take any spin $s$ of our choice, with an $(2s+1)^{L-1}$ equal number of degenerate and distinct eigenvalues ranging from $-\epsilon s$ to $\epsilon s$ in steps of $\epsilon$, where $\epsilon=2$ for half-integrer spins and $\epsilon=1$ for integrer spins, as introduced in (46)-(47). It is worth noting that alternative constructions of the higher-spin operators are possible. These constructions may lead to different commutation relations, such as: $[X_{j}^{(s)},Y_{k}^{(s)}]=in_{s}\delta_{jk}\,Z_{j}^{(s)}$, where $n_{s}=1/s$ with the additive relations $X_{i}^{(s)}X_{i}^{(s)}+Y_{i}^{(s)}Y_{i}^{(s)}+Z_{i}^{(s)}Z_{i}^{(s)}=\big{(}\frac{s+1}{s}\big{)}\,I_{N}$, as discussed in Refs. Xu:2019lhc ; Bhattacharjee:2022vlt . Another possible formulation is: $[X_{j}^{(s)},Y_{k}^{(s)}]=i\delta_{jk}\,Z_{j}^{(s)}$ with $X_{i}^{(s)}X_{i}^{(s)}+Y_{i}^{(s)}Y_{i}^{(s)}+Z_{i}^{(s)}Z_{i}^{(s)}=s(s+1)\,I_{N}$, as explored in Ref. Yin:2020oze . Despite these differing constructions, it is straightforward to demonstrate that they are related through appropriate scalings of the spin operators. For instance: $X_{j}^{(s)}\rightarrow\sqrt{\frac{s+1}{3s}}X_{j}^{(s)}$ (similar for $Y_{j}^{(s)}$ and $Z_{j}^{(s)}$) for the first case Xu:2019lhc ; Bhattacharjee:2022vlt and $X_{j}^{(s)}\rightarrow\sqrt{\frac{s(s+1)}{3}}X_{j}^{(s)}$ (similar for $Y_{j}^{(s)}$ and $Z_{j}^{(s)}$) for the second case Yin:2020oze . These scalings effectively normalize the eigenvalues to align with the conventions outlined in Table 1. Consequently, regardless of the specific construction chosen, the resulting operator statistics remain consistent and yield equivalent results.
@@ 6.2 Operator statistics at different times
Figure 8 illustrates the eigenvalue statistics of the time-evolved $s=1/2$ operator across three distinct phases of the system: the ergodic phase $\gamma=0$, the fractal phase $\gamma=1.5$, and the localized phase $\gamma=5$. These specific values of $\gamma$ are selected to ensure consistency with the eigenvalue statistics derived for the respective dimensions of the Hamiltonian; see Fig. 2.
At initial time $t=0$, the operator statistics are independent of the Hamiltonian and thus follow the multinomial distribution, depending on their number of distinct eigenvalues (see Table 1). This is also known as classical convolution. In the ergodic phase, the eigenvalue statistics of the operator exhibit convergence toward the arcsine distribution at the chosen timestamp, as predicted by free probability theory Camargo:2025zxr . This behavior reflects the expected random matrix-like dynamics within the ergodic regime. In contrast, the eigenvalue distributions in the fractal and localized phases deviate significantly from the arcsine distribution at the same timestamp. These deviations highlight the role of different nature of the quantum dynamics in the fractal and localized regimes, and we can speculate that quantum dynamics in these regimes of the Hamiltonian do not drive the operators (here $S_{1}(0)$ and $S_{L}(0)$) towards freeness (see also the discussion on the late time behavior of these statistics below).
Figure 9: The statistics of the eigenvalues of the time-evolved $s=1$ operator at the initial time $t=0$ (top, left) and late times $t=50$ in the ergodic (top, right), fractal (bottom, left), and localized (bottom, right) regimes. We choose $L=6$ (i.e., dimension of the Hamiltonian $N=(2s+1)^{L}=729$) and average over $1000$ independent realizations. The solid blue lines describe the distribution predicted by the free probability Eq. (14), i.e., Fig. 1 (b).
Figure 10: The statistics of the eigenvalues of the time-evolved $s=3/2$ operator at the initial time $t=0$ (top, left) and late times $t=50$ in the ergodic (top, right), fractal (bottom, left), and localized (bottom, right) regimes. We choose $L=5$ (i.e., dimension of the Hamiltonian $N=(2s+1)^{L}=1024$) and average over $500$ independent realizations of the Hamiltonian. The solid blue lines describe the distribution predicted by the free probability, i.e., Fig. 1 (c).
The same calculation is extended to spin $s=1$ and spin $s=3/2$ operators, as given in (46)-(47). The resulting operator statistics are illustrated in Fig. 9 and Fig. 10, corresponding to $s=1$ and $s=3/2$ respectively. The analytical expression for the distribution predicted by the free probability theory for $s=1$ is given in Eq. (14), derived previously in Camargo:2025zxr , while for $s=3/2$ the exact analytical expression of the free probability prediction is not known. The numerical findings for $s=1$ and $s=3/2$ remain consistent with the conclusions derived for the $s=1/2$.
@@
Figure 11: The statistics of the eigenvalues of the time-evolved $s=1/2$ (top panel), $s=1$ (middle panel), and $s=3/2$ (bottom panel) operator at the very late times $t\sim O(10^{4})$ in the fractal (left panels) and the localized (right panels) phases. The solid lines in all panels describe the distribution predicted by the free probability. The system parameters are the same as Fig. 8 – Fig. 10. In the fractal regime, the operator statistics converge to a specific distribution that deviates from the free probability prediction. Distinct spikes are observed at the same points as those of at $t=0$, indicating that the time-evolved operator retain some ‘memory’ of the initial distribution. Conversely, in the localized regime, the operator statistics remain markedly different from the results predicted by free probability, indicating the persistence of strong localization effects.

### v3

@@ 5.2 Correlation functions in the ergodic phase from asymptotic freeness
where $R$ is the radius of the semicircle distribution. For the RP Hamiltonian, in the regime of large $N$ and small $\gamma$, the radius scales approximately as $R=2/N^{\gamma/2}$, consistent with the structure of the Hamiltonian. [^14] For finite $N$, the semicircle formula describes the DOS up to some $\gamma\leq\gamma_{*}$ with $\gamma_{*}\geq 0$ in the ergodic phase. For example, with $N=500$, we observed that the Wigner semicircle describes the DOS well as long as $\gamma\leq 0.5$.
Figure 5: Two-point function $F_{2}^{(\gamma)}(t)$ (left) and four-point OTOC $F_{4}^{(\gamma)}(t)$ (right), shown in the ergodic phase (top row) and in the non-ergodic phases (bottom row). Parameters are set to $i=j=1$, $L=8$, and results are averaged over 100 realizations of the Hamiltonian.
For even values of $n$, the moments of the Hamiltonian obtained from the above DOS are given by $\langle H^{n}\rangle=C_{n}(R/2)^{2n}$, while for odd $n$, the moments vanish. Here, $C_{n}$ are Catalan numbers. Using (52), the two-point function can be expressed as
$\langle S_{i}(0)S_{j}(t)\rangle=\langle{S_{i}S_{j}}\rangle~\sum_{k,l}\frac{(it)^{2k}(-it)^{2l}}{k!~l!}C_{2k}C_{2l}\Big(\frac{R}{2}\Big)^{2(k+l)}=\langle{S_{i}S_{j}}\rangle\left(\frac{2J_{1}(Rt)}{Rt}\right)^{2}\,.$ (54)
Since the ensemble-averaged analytically continued partition function at infinite temperature is
$\langle Z(t)\rangle=\int\textrm{d}E\,\rho(E)~e^{-iEt}~=\frac{2J_{1}(Rt)}{Rt}\,.$ (55)
The time-dependent term on the right-hand side of the final expression in (54) corresponds to the part of the spectral form factor arising from the disconnected contribution of the two-point correlation function. For $\gamma=0$, the expression for $\langle Z(t)\rangle$ reduces to the Gaussian (e.g., GOE) result: $J_{1}(2t)/t$, as expected.
Next, we consider the four-point OTOC. For simplicity, we focus on the case of initially disjoint operators with zero one-point functions. This correlation function can be written as
$\langle S_{i}(0)S_{j}(t)S_{i}(0)S_{j}(t)\rangle=\sum_{k,l,m,n}\frac{(it)^{k+m}(-it)^{l+n}}{k!~l!~m!~n!}\langle{S_{i}H^{k}S_{j}H^{l}S_{i}H^{m}S_{j}H^{n}}\rangle\,.$ (56)
Since the initial operators are assumed to be non-overlapping, only a single diagram gives a non-zero contribution in the moment expansion in the large-$N$ limit, due to the freeness between the Hamiltonian $H$ and the operators $S_{i}$ (and $S_{j}$) in the ergodic phase of the RP model. Proceeding similarly to the two-point function, this contribution can be evaluated as Camargo:2025zxr :
$\langle S_{i}(0)S_{j}(t)S_{i}(0)S_{j}(t)\rangle=\langle{S_{i}S_{j}S_{i}S_{j}}\rangle\left(\frac{2J_{1}(Rt)}{Rt}\right)^{4}~,~~i\neq j~.$ (57)
Once again, the final expression corresponds to the disconnected contribution of the spectral correlation function to the four-point spectral form factor Cotler:2017jue .
Figure 6: Behavior of the $2n$-point function for $n=1$ to $n=5$ in the ergodic phase ($\gamma=0$), governed by a fully GOE Hamiltonian. The black dashed line is the analytic result (58) with $R=2$ (i.e., $\gamma=0$). Here, we set $i=j=1$, $L=8$, and average over 100 realizations of the Hamiltonian.
Simplifications for the spin-1/2 case
For $s=1/2$, expressions (54) and (56) can be further simplified using $\langle S_{i}S_{j}\rangle=\delta_{ij}$ and $\langle S_{i}S_{j}S_{i}S_{j}\rangle=\langle S_{i}^{2}S_{j}^{2}\rangle=1$ for $i\neq j$, yielding the simple results:
$F_{2}^{(\gamma)}(t)$ $=\langle S_{i}(0)S_{j}(t)\rangle=\delta_{ij}\,\left(\frac{2J_{1}(Rt)}{Rt}\right)^{2}\,$, (58)
$F_{4}^{(\gamma)}(t)$ $=\langle S_{i}(0)S_{j}(t)S_{i}(0)S_{j}(t)\rangle=\left(\frac{2J_{1}(Rt)}{Rt}\right)^{4}\,,\quad i\neq j\,.$ (59)
We remind the reader that the expressions for the correlation functions derived above using free probability theory are valid strictly in the $N\rightarrow\infty$ limit. For finite $N$, numerical results may exhibit deviations from these predictions; however, these differences vanish as $N$ becomes large Camargo:2025zxr .
Figure 5 shows the behavior of the two-point function $F_{2}^{(\gamma)}(t)$ (Eq.(51) with $n=1$ and $i=j=1$) and the four-point OTOC $F_{4}^{(\gamma)}(t)$ (Eq.(51) with $n=2$ and $i=j=1$) across the ergodic and non-ergodic phases for $s=1/2$. In the top row, both $F_{2}^{(\gamma)}(t)$ and $F_{4}^{(\gamma)}(t)$ decay to zero in the ergodic phase, with the decay becoming progressively slower as $\gamma$ increases. In contrast, the bottom row shows that for $\gamma>1$, the correlation functions no longer decay to zero. These observations support the picture that, in the ergodic phase, the algebra of observables at $t=0$ becomes asymptotically free from the algebra of observables at later times, with the apparent timescale for the onset of freeness or free time (to be precisely defined later) growing with $\gamma$. By contrast, in the non-ergodic phases, this mechanism fails, as $F_{2}^{(\gamma)}(t)$ and $F_{4}^{(\gamma)}(t)$ saturate to nonzero values.
To further investigate the emergence of asymptotic freeness in the ergodic phase, it is necessary to check whether higher-order OTOCs also decay to zero. As illustrated in Fig. 6, for a fixed $\gamma$, higher-order OTOCs indeed decay to zero, with the decay becoming increasingly rapid for larger values of $n$.
Finally, in Fig. 7, we compare the numerical results for $F_{2}^{(\gamma)}(t)$ and $F_{4}^{(\gamma)}(t)$ at small values of $\gamma$ and $s=1/2$ with the analytical expressions (58) and (59). The analytical formulas accurately reproduce the data at early times and deviate only when the correlation functions approach very small values. This discrepancy arises because we are working with $L=8$, corresponding to a Hilbert space dimension $N=(2s+1)^{L}=256$. For such a relatively small Hilbert space, the correlation functions do not decay exactly to zero but instead exhibit residual fluctuations, resulting from the finite-size effects, that decrease as $L$ increases. By contrast, our analytical formulas are strictly valid in the limit $N\to\infty$. The resulting mismatch, though small, is visible in the log-scale plots, where deviations are more easily amplified.
Figure 7: Log-scale comparison of the numerical and analytical results of the two-point functions $F_{2}^{(\gamma)}(t)$ with $i=j=1$ (left) and four-point OTOCs $F_{4}^{(\gamma)}(t)$ with $i=1$, $j=L$ (right). Each panel includes three curves for $\gamma=0$ (green), $\gamma=0.1$ (blue), and $\gamma=0.2$ (orange). Dots indicate numerical results, while solid lines represent the analytic predictions from Eqs. (58) (left panel) and (59) (right panel). Here, we set $L=8$ and average over 100 realizations of the Hamiltonian.
@@ 5.3 Cumulative OTOCs
Figures 8 and 9 illustrate the behaviour of the cumulative partial sum $\mathcal{F}_{p}^{(\gamma)}(t):=\mathcal{F}_{p}^{(\gamma)}(f\,;t)$ for the choice of weights $f(n)=1/n$ and $f(n)=1/n^{2}$ respectively in the ergodic (left) and fractal (right) regimes, for $s=1/2$ and $L=8$. Since higher-point functions decay faster than lower-point functions Cotler:2017jue , in practice, only a finite $p$ suffices to produce reasonable results.
@@
$\sum_{n=1}^{\infty}\frac{1}{n^{k}}\,e^{-2n\chi t}=\textrm{Li}_{k}(e^{-2\chi t})\approx e^{-2\chi t}\,,\quad(\mathrm{large}~t)\,.$ (64)
This is indeed representative of the numerical results for the ergodic phase (left) in Figs. 8 and 9, where the horizontal red dashed line in the latter case corresponds precisely to $\zeta(2)=\pi^{2}/6$. The exponential decay is represented by the black dashed line in Figs. 8 and 9 (left), demonstrating the eventual vanishing of the cumulative OTOC. The $\gamma=0$ result primarily shows the result when the Hamiltonian is strictly controlled by the GOE matrices in the RP Hamiltonian. The same conclusion holds for other classes of Gaussian ensembles.
@@
In the fractal phase, the cumulative sum undergoes a decay, albeit at a much slower rate compared to the ergodic regime. As illustrated in the inset of Figs. 8 and 9 (right), at late times, the cumulative OTOC does not approach zero; rather, it decays to a small, non-zero value. This phenomenon reflects a form of “memory” retained within the operator statistics, which we will see in the subsequent section. In contrast, within the localized phase, the cumulative partial sum exhibits minimal decay, remaining almost constant at its initial value.
6 Eigenvalue statistics of sum of operators from the free probability theory
In this section, we investigate the eigenvalue statistics of the sum of two operators, one of which is time-evolved by a Hamiltonian drawn from the RP ensemble, and compare it with the prediction of free probability theory discussed in Sec. 2.1. Specifically, our focus lies on the eigenvalue statistics of the following operator:
@@ 6 Eigenvalue statistics of sum of operators from the free probability theory
for specific values of the time parameter when the temporal evolution is generated by the RP model Hamiltonian, as defined in (20). Here $S_{1}(0)$ and $S_{L}(0)$ represent the spin operators at the initial time $t=0$, defined in (47) and (48). The total sum of the operators $S(t)$ constructed through this formalism, encapsulates the effect of the RP Hamiltonian (20) over time. Our primary objective is to systematically analyze the eigenvalue statistics of $S(t)$ at various time instants $t$. By varying $\gamma$, i.e., the parametric regime of the RP Hamiltonian, distinct phases (ergodic, fractal, or localized) or structures in the operator statistics are anticipated, potentially exhibiting novel characteristics inspired by predictions from free probability theory.
@@ 6.1 Choice of the initial operators
In the following analysis, we consider the operators $S_{1}(0)$ and $S_{L}(0)$ to be spin operators with spin $s$. This formulation generalizes the construction and enables comparisons against free probability predictions for various spin values, as obtained through analytical methods outlined in Section 2. The operators $S_{1}(0)$ and $S_{L}(0)$ are defined as tensor products of spin operators as in (47)-(48).
Figure 10: The statistics of the eigenvalues of the operator (65) with $s=1/2$ at the initial time $t=0$ (top, left) and late times $t=50$ in the ergodic (top, right), fractal (bottom, left), and localized (bottom, right) regimes. We choose $L=8$ (i.e., dimension of the Hamiltonian $N=(2s+1)^{L}=256$) with $5000$ ensemble averages. The solid blue lines describe the arsine distribution (12) i.e., Fig. 1 (a).
@@ 1 $S_{1},S_{L}$ $\{1,0,-1\}$
This is a convention that we follow throughout our discussion. More generally, we take any spin $s$ of our choice, with an $(2s+1)^{L-1}$ equal number of degenerate and distinct eigenvalues ranging from $\epsilon s$ to $\epsilon s$ in steps of $\epsilon$, where $\epsilon=2$ for half-integrer spins and $\epsilon=1$ for integrer spins, as introduced in (47)-(48). It is worth noting that alternative constructions of the higher-spin operators are possible. These constructions may lead to different commutation relations, such as: $X_{j}^{(s)},Y_{k}^{(s)}]=in_{s}\delta_{jk}\,Z_{j}^{(s)}$, where $n_{s}=1/s$ with the additive relations $X_{i}^{(s)}X_{i}^{(s)}+Y_{i}^{(s)}Y_{i}^{(s)}+Z_{i}^{(s)}Z_{i}^{(s)}=\big(\frac{s+1}{s}\big)\,I_{N}$, as discussed in Refs. Xu:2019lhc ; Bhattacharjee:2022vlt . Another possible formulation is: $X_{j}^{(s)},Y_{k}^{(s)}]=i\delta_{jk}\,Z_{j}^{(s)}$ with $X_{i}^{(s)}X_{i}^{(s)}+Y_{i}^{(s)}Y_{i}^{(s)}+Z_{i}^{(s)}Z_{i}^{(s)}=s(s+1)\,I_{N}$, as explored in Ref. Yin:2020oze . Despite these differing constructions, it is straightforward to demonstrate that they are related through appropriate scalings of the spin operators. For instance: $X_{j}^{(s)}\rightarrow\sqrt{\frac{s+1}{3s}}X_{j}^{(s)}$ (similar for $Y_{j}^{(s)}$ and $Z_{j}^{(s)}$) for the first case Xu:2019lhc ; Bhattacharjee:2022vlt and $X_{j}^{(s)}\rightarrow\sqrt{\frac{s(s+1)}{3}}X_{j}^{(s)}$ (similar for $Y_{j}^{(s)}$ and $Z_{j}^{(s)}$) for the second case Yin:2020oze . These scalings effectively normalize the eigenvalues to align with the conventions outlined in Table 1. Consequently, regardless of the specific construction chosen, the resulting operator statistics remain consistent and yield equivalent results.
@@ 6.2 Operator statistics at different times
Figure 10 illustrates the eigenvalue statistics of the operator (65) with $s=1/2$ across three distinct phases of the system: the ergodic phase $\gamma=0$, the fractal phase $\gamma=1.5$, and the localized phase $\gamma=5$. These specific values of $\gamma$ are selected to ensure consistency with the eigenvalue statistics derived for the respective dimensions of the Hamiltonian; see Fig. 2.
At the initial time $t=0$, the operator statistics are independent of the Hamiltonian and thus follow the multinomial distribution, depending on their number of distinct eigenvalues (see Table 1). This is what is expected from the classical convolution of the distributions of the eigenvalues of the individual operators. In the ergodic phase, the eigenvalue statistics of the operator exhibit convergence toward the arcsine distribution at the chosen timestamp, as predicted by free probability theory Camargo:2025zxr . This behavior reflects the expected random matrix-like dynamics within the ergodic regime. In contrast, the eigenvalue distributions in the fractal and localized phases deviate significantly from the arcsine distribution at the same timestamp. These deviations highlight the role of different nature of the quantum dynamics in the fractal and localized regimes, and we can speculate that quantum dynamics in these regimes of the Hamiltonian do not drive the operators (here $S_{1}(0)$ and $S_{L}(0)$) towards freeness (see also the discussion on the late time behavior of these statistics below).
Figure 11: The statistics of the eigenvalues of the operator (65) with $s=1$ at the initial time $t=0$ (top, left) and late times $t=50$ in the ergodic (top, right), fractal (bottom, left), and localized (bottom, right) regimes. We choose $L=6$ (i.e., dimension of the Hamiltonian $N=(2s+1)^{L}=729$) and average over $1000$ independent realizations. The solid blue lines describe the distribution predicted by the free probability Eq. (14), i.e., Fig. 1 (b).
Figure 12: The statistics of the eigenvalues of the operator (65) with $s=3/2$ at the initial time $t=0$ (top, left) and late times $t=50$ in the ergodic (top, right), fractal (bottom, left), and localized (bottom, right) regimes. We choose $L=5$ (i.e., dimension of the Hamiltonian $N=(2s+1)^{L}=1024$) and average over $500$ independent realizations of the Hamiltonian. The solid blue lines describe the distribution predicted by the free probability, i.e., Fig. 1 (c).
The same computations can be extended to spin $s=1$ and spin $s=3/2$ operators given in (47)-(48). The resulting eigenvalue statistics are illustrated in Fig. 11 and Fig. 12, corresponding to $s=1$ and $s=3/2$ respectively. The analytical expression for the distribution predicted by the free probability theory for $s=1$ is given in Eq. (14), derived previously in Camargo:2025zxr , while for $s=3/2$ the exact analytical expression of the free probability prediction is not known. The numerical findings for $s=1$ and $s=3/2$ remain consistent with the conclusions derived for the $s=1/2$.
@@
Figure 13: The statistics of the eigenvalues of the operator (65) with $s=1/2$ (top panel), $s=1$ (middle panel), and $s=3/2$ (bottom panel) at the very late times $t\sim O(10^{4})$ in the fractal (left panels) and the localized (right panels) phases. The solid lines in all panels describe the distribution predicted by the free probability. The system parameters are the same as Fig. 10 – Fig. 12. In the fractal regime, the operator statistics converge to a specific distribution that deviates from the free probability prediction. Distinct spikes are observed at the same points as those of at $t=0$, indicating that the time-evolved operator retain some ‘memory’ of the initial distribution. Conversely, in the localized regime, the operator statistics remain markedly different from the results predicted by free probability, indicating the persistence of strong localization effects.